# Displacement Memory Effect from Supersymmetry

### v1

@@ Abstract
We explain the recent results regarding the displacement memory effect (DME) of plane gravitational waves using supersymmetric quantum mechanics. This novel approach stems from the fact that both geodesic and Schrödinger equations are Sturm-Liouville boundary value problems. Supersymmetry provides a unified framework for Pöschl-Teller and Scarf profiles and restores the critical values of wave amplitudes for DME in a natural way. Within our framework, we obtain a compact formula for DME in terms of the asymptotic values of the superpotential and the geodesics. In addition, this new technique enables us to build plane and gravitational waves with 2-transverse directions using superpartner potentials. Lastly, we study DME within a singular wave profile inspired by supersymmetric quantum mechanics, which shows the broader applicability of our method.
@@ 1 Introduction
The memory effect of plane waves can be mainly divided into two: the displacement memory effect (DME) [1] and the velocity memory effect (VME) [2, 3, 4]. These effects are of interest as they can be captured by future gravitational wave detectors. In this paper, we address the problem of DME of plane and gravitational waves (GWs) with a novel perspective using supersymmetric quantum mechanics (SUSY QM) [5, 6]. Our main motivation is to investigate the recent results showing the existence of DME in Pöschl-Teller and Scarf GW profiles in the light of SUSY QM.
@@
Previous studies (see [7, 8] and references therein) suggested that plane waves might only have a VME: after the passing of a short burst of a gravitational wave, transverse geodesics have a non-vanishing constant velocity. Only very recently, it has been reported that plane GWs, under special conditions, also can exhibit a DME: permanent displacement with vanishing velocity. Namely, DME occurs only in particular setups with fine-tuned parameters [9, 10, 11] (see also [12]).
@@
Our starting point is the fact that geodesics having DME conditions do satisfy a Sturm-Liouville problem with Neumann boundary conditions. This leads to an equation that we are already familiar with from non-relativistic QM. Therefore, DME of plane GWs and the related conditions can be sought for by using the Schrödinger theory, and the desired geodesics, in principle, could be found analytically by the methods of SUSY QM [5, 6].
@@
This common mathematical origin of geodesics of plane waves and quantum mechanics sheds light on the recent results of those half-waves and magical numbers of the former by relating it to the bound states and quantum numbers of the latter. We show that SUSY elegantly and simply produces these particular setups. We also borrow a singular potential from SUSY QM and discuss the possibility of finding DME within singular GWs, such as the Lukash GW studied in cosmological models.
@@
Plane waves, being models for gravitational radiation, are already significant spacetimes on their own [13]. They are Penrose limits of any spacetime [14], and lower dimensional mechanics can be embedded into them via the Eisenhart lift [15, 16, 17].
@@
As a novelty, we now show that the necessary conditions for their DME relate them to integrable potentials in SUSY QM. Using this interrelation, we obtain the geodesics with DME out of SUSY methods in a compact and beautiful manner. SUSY also helps us to understand the conditions to build exact GWs exhibiting DME in both directions.
@@
Let us note the work of C. Duval et al. [18], who discussed Schrödinger theory in the Bargmann framework. The geodesic equations of spinning particles in gravitational backgrounds were worked out using supersymmetry in [19]. Our study differs from those both in its content and in its scope.
@@
The outline of this paper is the following: In Section 2, we introduce our main motivation, which is DME of the plane waves. Then, we apply SUSY methods of quantum mechanics to plane wave geodesics and discuss how DME occurs in Section 3. 4-dimensional plane waves and their DME are discussed in Section 4. In Section 5, we commented on the DME of singular wave profiles. In our last Section, Section 6, we summarize our results, add our comments, and ask further questions. For completeness, we provide the necessary tools of SUSY QM in the Appendix A.
@@ 2 Plane waves and memory effects
where $U,V$ are light-cone coordinates and $X^{i}$ with $i=1,...,d$, are transverse coordinates.
To discuss DME and VME, we consider geodesic motion. Plane waves (2.1) have a covariantly constant null Killing vector $\frac{\partial}{\partial V}$. Therefore, the coordinate $U$ can be chosen as an affine parameter, and geodesic equations for the remaining coordinates turn out to be
$\frac{d^{2}X^{i}}{dU^{2}}-K_{ij}X^{j}=0,$ (2.2a)
$\frac{d^{2}V}{dU^{2}}+\frac{1}{2}\frac{dK_{ij}}{dU}X^{i}X^{j}+2K_{ij}X^{i}\frac{dX^{i}}{dU}=0.$ (2.2b)
@@
$V(U)=-\frac{1}{2}X^{i}\frac{dX^{i}}{dU}+c_{1}U+c_{2}.$ (2.3)
The constants $c_{1}$ and $c_{2}$ are to be determined from the initial conditions. So, the most crucial part of the geodesic problem (thus, of the memory effect) is to solve (2.2a), which, when $K_{ij}$ depends on time, is a Sturm-Liouville problem.
Only a few exact solutions are known, and generally, we confine ourselves to numerical methods (see, e.g, [8]). Analytical or numerical, both types of studies show that geodesics of plane GWs have, generically, a non-vanishing constant velocity after the wave has passed, confirming the VME effect found in [2, 3, 4].
Therefore, it comes as a surprise that the Gaussian profile with just 1 transverse direction
@@
yields DME if certain critical values are chosen for the amplitude, $k=k_{crit}$ [9, 10]. This confirms the statement of Zeldevich and Polnarev [1], who advocated that the relative velocity will be zero and there will only be a pure displacement at distant times. In that manner, the geodesic should satisfy the Neumann boundary conditions
@@
and the related equation (2.2a) becomes a Sturm-Liouville (SL) boundary value problem.
In their quest for such magical values $k=k_{crit}$ authors of [9, 10] were motivated to replace Gaussian profile (2.4) with that of Pöschl-Teller (PT)
$K^{PT}(U)=\frac{k}{2\cosh^{2}U},$ (2.6)
which is a good approximation of the Gaussian (2.4).
While for $K^{G}(U)$ (and its derivatives) only numerical solutions are obtained, geodesics of PT profile (2.6) with the condition (2.5) can be found analytically: DME occurs for $k_{crit}=4m(m+1)$ where $m$ is a positive integer. Moreover, their very recent paper [11] reports that the scarf profile
$K^{scarf}(U)=-2g\frac{\sinh U}{\cosh^{2}U}.$ (2.7)
is also a good approximation to flyby [1] described by the derivative of (2.4). Similar to the PT profile (2.6), the Scarf profile is also capable of having a DME if magic parameter values are chosen. In that case, it is even possible to create an exact plane GW with 2-transverse directions, both yielding DME geodesics. This is true also for the derivative of the Gaussian profile.
Let us mention that a PT-type wave profile (2.6) was also employed in [20] to study gravitational wave memories (DME and VME) within Ellis-Bronnikov wormhole geometries. In [21], the same profile was used in Kundt wave spacetimes to indicate a relation between curvature and memory. The authors showed the existence of DME numerically. In [22], similar profiles are adopted to study the so-called ”B-memory” of vacuum plane waves.
Leaving those numerical studies aside, we will focus our attention on analytical/exact results for (2.6) and (2.7). We will treat the related geodesic equations (2.2a) as a non-relativistic Schrödinger equation. The latter is exactly solvable for a few cases. A list of them [5, 6] includes (2.6) and (2.7). Thus, the methods of supersymmetry are directly applicable to the related plane GW geodesics. In that way, we will reveal the origin of DME of plane waves in the context of SUSY QM, which restores the magical values of parameters and necessary conditions naturally. We believe that this similarity of plane wave geodesics with supersymmetry in quantum mechanics is interesting in its own right.
@@ 3 Creating plane and gravitational waves via SUSY QM
The transverse part of the geodesic equations in plane waves (2.2a) together with (2.5) constitute the SL boundary value problem [^1] For simplicity, we consider uncoupled equations with vanishing off-diagonal elements.. Therefore, an exactly solvable geodesic equation of this form is expected to have a correspondence in SUSY QM where exact solutions can be derived with its own elegant methods.
Here, we will explicitly reveal the correspondence between geodesics of plane waves in Section 2 and quantum mechanics by deriving the results of [9, 10, 11] with the methods of SUSY. Basic tools of SUSY QM can be recalled from our Appendix A whenever necessary.
@@
    The usage superpotential $W(U)$ (A.1) allows us to unify the recent works about DME of plane GWs in an elegant way.
@@
    Incorporation of supersymmetric methods explain the magical amplitude values and half waves of DME effect.
@@
Let us achieve our first goal.
@@ 3.1 Unified wave profiles via SUSY QM
Consider the superpotential
$W(U)=A\text{tanh}(\alpha U)+B\text{sech}(\alpha U).$ (3.1)
This potential is noted in the list of shape invariant potentials in [5, 6] (for earlier studies, see [23, 24]). $A,B,\alpha$ are free parameters for now. Out of this superpotential (3.1), one can derive 2 partner potentials (A.2)
$V_{1}$ $=W^{2}-\frac{dW}{dU}=(B^{2}-A^{2}-A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A+\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)+A^{2}\qquad\qquad$ (3.2a)
$V_{2}$ $=W^{2}+\frac{dW}{dU}=(B^{2}-A^{2}+A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A-\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)+A^{2}.\qquad\qquad$ (3.2b)
$V_{1}$ and $V_{2}$ are related with the shape invariance condition (A.5) with $A$ being the shape invariance parameter
$V_{2}(U;A)=V_{1}(U;A-\alpha)+\alpha(2A-\alpha).$ (3.3)
Using the condition (3.3) and shifting $A$, exact bound state spectrum for $V_{1}$ (hence $V_{2}$) can be found
$E_{n}=A^{2}-(A-n\alpha)^{2},\quad n=0,1,2,...$ (3.4)
Above $n=0$ is the ground state with $E_{0}=0$, $n=1$ is the first excited bound state, etc. Please note that their wavefunctions vanish at $U\to\pm\infty$. Below, we will see that their first derivatives are also vanishing at distant times so that they satisfy DME conditions (2.5).
We put the associated Schrödinger equation (A.3) into the form of a geodesic equation (2.2a)
$\frac{d^{2}\psi_{n}}{dU^{2}}-\left[(A-n\alpha)^{2}+(B^{2}-A^{2}-A\alpha)\ \text{sech}^{2}(\alpha U)+B(2A+\alpha)\ \text{sech}(\alpha U)\text{tanh}(\alpha U)\right]\psi_{n}=0.$ (3.5)
Up to a normalization, the ground state wave function can be found (A.4)
$\psi_{0}(U)=e^{-\frac{1}{\alpha}\left(A\ln\big{(}\cosh(\alpha U)\big{)}+B\tan^{-1}\big{(}\sinh(\alpha U)\big{)}\right)}.$ (3.6)
In order to pass to the geodesics (2.2), we make the replacement $\psi(U)\to X(U)$ and set $\alpha=1$ in the above expressions. It will be immediately recognized that (3.5) is composed of 2 parts: PT profile (2.6) in [9, 10] and Scarf profile (2.7) in [11]. Therefore, we have achieved our first goal of a unified framework. It is precisely this unified view that will allow us to understand DME of the related profiles in a systematic way, explaining the quantized values of magical amplitudes.

### v3

@@ Abstract
We explain the recent results on the displacement memory effect (DME) of plane gravitational waves using supersymmetric quantum mechanics. This novel approach stems from that both the geodesic and the Schrödinger equations are Sturm-Liouville boundary value problems. Supersymmetry provides a unified framework for the Pöschl-Teller and the Scarf profiles and yields the critical values of the associated wave amplitudes for DME in a natural way. Within our framework, we obtain a compact formula for DME in terms of the asymptotic values of the superpotential and the geodesics. In addition, this new technique enables us to build plane and gravitational waves with 2-transverse directions using superpartner potentials. Lastly, we study DME within a singular wave profile inspired by supersymmetric quantum mechanics, which demonstrates the broader applicability of our method.
@@ 1 Introduction
The memory effect of plane gravitational waves can be divided into two types: the displacement memory effect (DME) [1] and the velocity memory effect (VME) [2, 3, 4]. These effects are of interest as they can hopefully be captured by future gravitational wave detectors [5, 6]. In this paper, we focus our attention at the problem of DME of the plane gravitational waves (GWs) with a novel perspective by applying techniques borrowed from supersymmetric quantum mechanics (SUSY QM) [7, 8, 9, 10, 11].
@@
Previous studies (see [12, 13] and references therein) suggested that plane waves might only have a VME: after the passing of a short burst of a gravitational wave, transverse geodesics have a non-vanishing constant velocity. Only very recently it has been reported that plane GWs, under special conditions, can also exhibit a DME, that is, a permanent displacement with vanishing velocity after the wave has passed. According to [14, 15, 16] (see also [17]), DME occurs only in particular setups with fine-tuned parameters. Our main motivation is to investigate these recent results about the DME for the Pöschl-Teller (PT) (2.6) and the Scarf GW (2.7) profiles from the point of view of SUSY QM. We show that there is a hidden supersymmetry behind the DME phenomena of the PT and the Scarf GWs which unifies those profiles and determines the critical wave amplitudes in an elegant manner.
@@
Our starting point is that the geodesics with DME conditions satisfy a Sturm-Liouville (SL) problem with Neumann boundary conditions. We are already familiar with such equations from another context, namely non-relativistic QM. Therefore, the DME conditions of the plane GWs can be sought for by using the Schrödinger theory. The desired geodesics could, in principle, be found analytically by the methods of SUSY QM [7, 8, 9, 10].
@@
This common mathematical origin of geodesics of plane waves and quantum mechanics sheds light on the recent results of those half-waves and magical numbers of the former [14, 15, 16] by relating them to the bound states and the quantum numbers of the latter. Namely, we begin with a superpotential (3.9) and create partner potentials (3.10) including both the Pöschl-Teller and the Scarf profiles at the same time. Then, SUSY leads to those particular setups and solutions in a natural way. We also borrow a singular potential from SUSY QM and discuss the possibility of finding DME within singular GWs.
@@
Plane waves, being models for gravitational radiation, are already significant spacetimes on their own [18]. They are Penrose limits of any spacetime [19], and lower dimensional mechanics can be embedded into them via the Eisenhart-Duval lift [20, 21, 22].
@@
We show, for the first time we believe, that the necessary conditions for DME relate plane waves to integrable potentials in SUSY QM. Using this novel interrelation, we obtain the geodesics with DME by using SUSY methods in a compact and beautiful manner. SUSY also helps us to understand the conditions to build exact GWs exhibiting DME in both directions.
@@
Let us note that the geodesic equations of spinning particles in gravitational backgrounds were worked out using supersymmetry in [23]. Our study differs from that both in its content and in its scope.
@@
The outline of the paper is the following: In Section 2, we introduce our main motivation, which is the DME of plane waves. Then, in Section 3, we apply SUSY methods of quantum mechanics to plane wave geodesics and discuss how DME occurs. 4-dimensional plane waves and their DME are discussed in Section 4. In Section 5, we comment on the DME for singular wave profiles. In the last section, Section 6, we summarize our results, add our comments, and ask further questions.
@@ 2 Plane waves and memory effects
where $U,V$ are light-cone coordinates and $X^{i}$ with $i=1,...,d$, are transverse coordinates. The profile $K_{ij}(U)$ is a symmetric $d\times d$ matrix. The plane wave (2.1) is known to be an exact plane GW if its profile is traceless.
To discuss DME and VME, we consider the geodesic motion. Plane waves (2.1) have a covariantly constant null Killing vector $\frac{\partial}{\partial V}$. Thus, the coordinate $U$ can be chosen as an affine parameter, and geodesic equations for the remaining coordinates turn out to be
$\frac{d^{2}X^{i}}{dU^{2}}-K^{i}{}_{j}X^{j}=0,$ (2.2a)
$\frac{d^{2}V}{dU^{2}}+\frac{1}{2}\frac{dK_{ij}}{dU}X^{i}X^{j}+2K_{ij}X^{i}\frac{dX^{j}}{dU}=0.$ (2.2b)
@@
$V(U)=-\frac{1}{2}X^{i}\frac{dX_{i}}{dU}+c_{1}U+c_{2}.$ (2.3)
The constants $c_{1}$ and $c_{2}$ are to be determined from the initial conditions. So, the most crucial part of the geodesic problem (thus, of the memory effect) is to solve (2.2a), which is an SL problem when $K_{ij}$ depends on $U$.
Only a few exact solutions are known, and we usually confine ourselves to numerical methods (see, e.g, [13]). Analytical or numerical, both types of studies show that the geodesics of plane GWs have, generically, a non-vanishing constant velocity after the wave has passed, confirming the VME considered in [2, 3, 4]. Therefore, it comes as a surprise that the Gaussian profile with $d=1$
@@
yields DME if certain critical values are chosen for the amplitude, $k=k_{crit}$ [14, 15]. This confirms the statement of Zel’dovich and Polnarev [1], who advocated that the relative velocity will be zero and there will only be pure displacement at distant times. In that manner, the geodesic should satisfy the Neumann boundary conditions
@@
and the related equation (2.2a) becomes an SL boundary value problem.
In their quest for such magical values $k=k_{crit}$, authors of [14, 15] were motivated to replace the Gaussian profile (2.4) with that of the Pöschl-Teller (PT)
$K^{PT}(U)=-\frac{k}{4\cosh^{2}U}.$ (2.6)
This profile yields a good approximation of the Gaussian (2.4). While for $K^{G}(U)$ only numerical solutions are available, the geodesics of the PT profile (2.6) with (2.5) can be found analytically: DME occurs for $k_{crit}=4m(m+1)$ where $m$ is a positive integer.
Moreover, a very recent paper [16] reports that the Scarf profile
$K^{scarf}(U)=-2g\frac{\sinh U}{\cosh^{2}U},$ (2.7)
can be used as an approximation to the derivative of (2.4) proposed to describe flyby [24]. Akin to the PT profile (2.6), the Scarf profile is also capable of having a DME if magic parameter values are chosen. It is even possible to create an exact plane GW with $d=2$, with both components yielding DME geodesics. Similar results were found numerically for the derivative Gaussian profile [16].
Let us mention that a PT-type wave profile (2.6) was also employed in [25, 26] to study DME and VME within wormhole geometries. In [27], the same profile was used in Kundt wave spacetimes to indicate a relation between the curvature and the memory. The authors showed the existence of DME numerically. In [28], similar profiles are adopted to study the B-memory of the vacuum plane waves and in [29] DME was studied in asymptotic black hole spacetimes.
Leaving those numerical studies aside, we employ SUSY to obtain exact results both for the PT profile (2.6) and for the Scarf profile (2.7). The list of exactly solvable potentials in [8, 9, 10] includes them within a single superpotential. Thus, the methods of SUSY are directly applicable to their geodesic problem which can be treated as a non-relativistic Schrödinger equation. Using this relation, we reveal the origin of DME of plane waves in the context of SUSY QM, which restores the magical values of parameters and necessary conditions naturally. We believe that this interrelation of the DME problem of plane GWs with SUSY QM is interesting in its own right.
@@ 3 Creating plane and gravitational waves via SUSY QM
The transverse part of the geodesic equations in plane waves (2.2a), together with (2.5), constitute an SL boundary value problem [^1] For simplicity, the profile $K_{ij}(U)$ is taken to be diagonal.. Therefore, analytically solvable geodesic equations of this form are expected to correspond to exact solutions derived with elegant methods in SUSY QM.
Here, we will explicitly reveal the correspondence between geodesics of plane waves in Section 2 and quantum mechanics by deriving the results of [14, 15, 16] with the methods of SUSY.
@@
    Using a superpotential $W(U)$ (3.1) allows us to unify the recent works about DME of plane GWs in an elegant way.
@@
    Incorporation of supersymmetric methods explain the magical amplitude values and the half waves of DME for those particular GWs.
@@
3.1 SUSY QM: A brief review
Below, we provide a brief, self-contained summary of SUSY methods. The interested reader may consult [8, 9, 10] for more details.
We begin with recalling the main elements of SUSY QM, namely its superpotential $W(U)$ and ladder operators $L,L^{\dagger}$,
$L=\frac{d}{dU}+W(U),\quad L^{\dagger}=-\frac{d}{dU}+W(U).$ (3.1)
They factorize the Hamiltonians $H_{1}$ and $H_{2}$ as
$H_{1}$ $=L^{\dagger}L=-\frac{d^{2}}{dU^{2}}+V_{1},\quad\quad V_{1}=W^{2}-\frac{dW}{dU},$ (3.2a)
$H_{2}$ $=LL^{\dagger}=-\frac{d^{2}}{dU^{2}}+V_{2},\quad\quad V_{2}=W^{2}+\frac{dW}{dU}.$ (3.2b)
$V_{1}$ and $V_{2}$ (3.2) are partner potentials sharing the same spectrum, except for the zero-energy ground state of $V_{1}$. The time-independent Schrödinger equation for the first potential $V_{1}$ is
$\left(-\frac{d^{2}}{dU^{2}}+V_{1}\right)\psi_{n}=E_{n}\psi_{n},$ (3.3)
where $n=0,1,2,...$ indexes the bound state. Up to a normalization factor, the zero energy ground state $\psi_{0}(U)$ can be found as
[…]
If $V_{1}$ and $V_{2}$ are shape invariant [8] i.e.,
$V_{2}(U;A_{1})=V_{1}(U;A_{2})+f(A_{1}),$ (3.5)
where $A_{1,2}$ are parameters [^2] Recall that plane wave profiles (2.4), (2.6) and (2.7) also have parameters like $k,g,..$., energy eigenvalues and eigenstates of (3.2) can be computed easily. Both $f(A_{1})$ and $A_{2}(A_{1})$ are functions of the first parameter $A_{1}$. This crucial property of shape invariance of SUSY QM will allow us to derive DME of plane GWs.
Due to (3.5), discrete eigenstates are found one by one
$\psi_{n}(U;A_{1})=L^{\dagger}(U;A_{1})\psi_{n-1}(U;A_{2}).$ (3.6)
For instance, if $A_{1}$ and $A_{2}$ are related by a simple relation $A_{1}-A_{2}=1$, then the excited state wave functions can be computed with the subsequent action of $L^{\dagger}$ (3.1)
$\psi_{n}(U;A)=L^{\dagger}(U;A)L^{\dagger}(U;A-1)...L^{\dagger}(U;A-n+1)\psi_{0}(U;A-n),$ (3.7)
where we set $A_{1}=A$. In that case, (3.6) simplifies as
$\psi_{n}(U;A)=\left(-\frac{d}{dU}+W(U;A)\right)\psi_{n-1}(U;A-1).$ (3.8)
For now, the parameter $A$ can take any value. However, when this framework is applied to the geodesics, the DME problem puts restrictions on such parameters.
@@ 3.2 Unified wave profiles via SUSY QM
Consider the following superpotential
$W(U)=A\,\tanh U+B\,\operatorname{sech}U,$ (3.9)
where $A,B$ are free parameters. It figures in the list of shape invariant potentials in [7, 8, 9, 10] [^3] For a historical remark, see [30].. Out of it, one can build two partner potentials (3.2)
$V_{1}$ $=(B^{2}-A^{2}-A)\ \operatorname{sech}^{2}U+B(2A+1)\ \operatorname{sech}U\tanh U+A^{2},\qquad\qquad$ (3.10a)
$V_{2}$ $=(B^{2}-A^{2}+A)\ \operatorname{sech}^{2}U+B(2A-1)\ \operatorname{sech}U\tanh U+A^{2}.\qquad\qquad$ (3.10b)
$V_{1}$ and $V_{2}$ are related via (3.5) where $A$ is the shape invariance parameter
$V_{2}(U;A)=V_{1}(U;A-1)+2A-1.$ (3.11)
Using (3.11) and shifting $A$, the exact bound state spectrum for $V_{1}$ (hence for $V_{2}$) can be found as [9]
$E_{n}=A^{2}-(A-n)^{2},\quad n=0,1,\cdots A.$ (3.12)
The ground state with $E_{0}=0$ corresponds to $n=0$, $n=1$ is the first excited bound state, etc. Note that $V_{1,2}$ vanish as $U\to\pm\infty$. Below, we will show that they admit wave functions with vanishing derivatives at infinities so that they satisfy the DME conditions (2.5).
We cast the associated Schrödinger equation (3.3) into the form of a geodesic equation (2.2a)
$\frac{d^{2}\psi_{n}}{dU^{2}}-\left[(A-n)^{2}+(B^{2}-A^{2}-A)\ \operatorname{sech}^{2}U+B(2A+1)\ \operatorname{sech}U\tanh U\right]\psi_{n}=0.$ (3.13)
Up to a normalization factor, the ground state wave function can be obtained from (3.4),
$\psi_{0}(U)=e^{-\left(A\ln(\cosh U)+B\tan^{-1}(\sinh U)\right)}.$ (3.14)
In order to pass to the geodesics (2.2), we make the replacement $\psi(U)\to X(U)$ in the above expressions. It will be immediately recognized that (3.13) is composed of 2 parts: the PT profile (2.6) in [14, 15] and the Scarf profile (2.7) in [16]. Therefore, we achieve our first goal of a unified framework. This unified view will allow us to understand DME of the related profiles in a systematic way by explaining the quantized values of the magical amplitudes.